Please square answers

A 1400-kg rocket has a net propulsion force of 30 kN (kiloNewtons). Over a short time period, it uniformly speeds up from an initial velocity of 30 m/s to a final velocity of 55 m/s. Assume that the mass of the rocket is constant during this time period and that the net force is along the direction of motion.

What is the change in speed of the rocket in m/s and in mph?
m/s
mph

Tries 0/2


What is the magnitude of the impulse delivered to the rocket during this time period?
kg*m/s

Tries 0/2


How long did it take for the rocket to speed up from 30 m/s to 55 m/s?
s

Tries 0/2


What is the magnitude of the acceleration of the rocket?
m/s2

Tries 0/2

What are the initial and final kinetic energies of the rocket in kilojoules (kJ)?
KE initial:  kJ
KE final:  kJ

Tries 0/2


What is the net work done on the rocket in kilojoules (kJ)?
kJ

Tries 0/2


How far did the rocket travel during this time period?
m
Answer #1

A)

Change in speed, v = (55 - 30) m/s x 2.24 = 56 mph

B)

Impulse = 1400 x (55 - 30) = 35000 kg m/s

C)

Acceleration, a = F/m = 30000/1400 = 21.43 m/s^2

Time, t = (55 - 30)/21.43 = 1.17 sec

D)

Acceleration, a = F/m = 30000/1400 = 21.43 m/s^2

E)

Initial energy, E1 = 0.5 x 1400 x 30^2 = 630 kJ

Final energy, E2 = 0.5 x 1400 x 55^2 = 2117.5 kJ

F)

Work = E2 - E1 = 1487.5 kJ

G)

Distance, d = 1487500/30000 = 49.58 m

Comment in case any doubt please rate my answer....